OpenStudy (anonymous):

When at the free-throw line for two shots, a basketball player makes at least one free throw 90% of the time. 80% of the time, the player makes the first shot, 80% of the time the player makes the second shot and 70% of the time she makes both shots. What is the conditional probability that the player makes the second shot given that she missed the first? Express your answer as a percentage to the nearest percentage point.

OpenStudy (kirbykirby):

Let \(F\) = player makes the first shot \(\overline{F}\) = player does not make the first shot \(S\) = player makes the second shot You are given: \( P(F \cup S)=0.9\) since at least one shot means they got First shot OR Second short or both (and there are only 2 shots in total) \(P(F)=0.8\) \(P(S)=0.8\) \(P(F \cap S)=0.7\) You are asked: to find \(P(S|\overline{F})\) \[ P(S|\overline{F})=\frac{P(S\cap\overline{F})}{P(\overline{F})}\] Notice that \(P(S)=P(S\cap F)+P(S\cap\overline{F})\\ \implies P(S\cap \overline{F})=P(S)-P(S\cap F)=0.8-0.7=0.1\) And \(P(\overline{F})=1-P(F)=1-0.8=0.2\) So: \[ P(S|\overline{F})=\frac{P(S\cap\overline{F})}{P(\overline{F})}=\frac{0.1}{0.2}=0.5=50\%\]

OpenStudy (anonymous):

Thank you very much.